OpenStudy (anonymous):

Given the function f(x) = x3 + x2 - 2x + 1, what is resulting function when f(x) is shifted to the left 1 unit?

hero (hero):

To shift \(f(x)\) \(b\) units to the left, evaluate \(f(x + b)\).

hero (hero):

In this case, \(b = 1\), so evaluate \(f(x + 1)\).

OpenStudy (anonymous):

Can you show me how to do that? I am very confused.

OpenStudy (anonymous):

@Hero

hero (hero):

\(f(x + 1) = (x + 1)^3 + (x + 1)^2 - 2(x + 1) + 1\)

hero (hero):

Expand the right side

OpenStudy (anonymous):

Ohh okay i see. Would I combine like terms now?

hero (hero):

There are no like terms to combine at the moment. You can't combine exponents like that. You have to expand each term first.

OpenStudy (anonymous):

Do I need to find the GCF? @hero

hero (hero):

Hang on a minute...

hero (hero):

So you could start by factoring out \(x + 1\) to get \((x + 1)((x + 1)^2 + (x + 1) - 2) + 1\) That would help make simplifying this much easier.

OpenStudy (anonymous):

And then what should I do after that?

hero (hero):

Expand \((x + 1)^2\)

OpenStudy (anonymous):

(x+1)(x+1) right?

hero (hero):

Correct \((x + 1)^2 = (x + 1)(x + 1)\)

OpenStudy (anonymous):

Alright, now what do I do?

hero (hero):

Multiply\( (x+1)(x+1)\) completely to quadratic form.

OpenStudy (anonymous):

x=-1

hero (hero):

\(\begin{align*}(x + 1)(x + 1) &= x(x + 1) + 1(x + 1) \\&= x^2 + x + x + 1 \\&=x^2 + 2x + 1\end{align*}\)

OpenStudy (anonymous):

Ohhh okay I understand now. I think I've got the answer to the whole problem, can you check it?

hero (hero):

Let's see what you have.

OpenStudy (anonymous):

f(x + 1) = x3 + 4x2 + 3x + 1

hero (hero):

Correct, but just remember to use a caret for exponents.

hero (hero):

Great job.

OpenStudy (anonymous):

Okay thank you!!

hero (hero):

Can you please show the work you did to arrive at that result?